OpenStudy (anonymous):

Need help with interest rate question then I am done! Please help! A sum of $15,000 is split, and the two parts are invested at 8% and 6%, respectively. If the interest for the 8% exceeds the interest from the 6% investment by $360, how much is invested at each rate?

OpenStudy (anonymous):

I really don't have a formula to solve this, but maybe if you guess and check you can figure this out. I'd start with splitting it evenly. 15000/2 = 7500 7500*0.08= 600 7500*0.06= 450 that's only a 150 difference, so you know you need more in the 8% account... try out a few numbers and I bet you'll find it! Also, sorry that I don't have a quick and easy formula for ya. I wish I could be of more help, but this is what I would do in your situation. :) good luck!

OpenStudy (anonymous):

Thanks for attempting no one else has even tried haha.

OpenStudy (anonymous):

I'd probably be able to think of something easier but I don't think I've had enough coffee yet :P

OpenStudy (johnweldon1993):

oh...this is just basically solving 1 equation and subbing in what you get here like this

OpenStudy (johnweldon1993):

So you have x + y = 15000 right? (x * .08) - (y * .06) = $360

OpenStudy (johnweldon1993):

so lets solve for x first y = 15000 - x right? Now plug that into the second equation (x * .08) - ((15000-x) * .06) = $360 now distribute that .08x - 900 + .06x = 360 combine like terms .14x = 1260 and divide by .14 x = 9000 Now sub back into the first to find y 9000 + y = 15000 y = 6000 To check (9000 * .08) - (6000 * .06) = $360 720 - 360 = 360? Yes it does :)